OpenStudy (anonymous):

what are the approximate solutions of the equation x^2+4x=-2 A (-4.45, .45) B (-3.41, .45) C (-.59, -4.45) D (-.59, -3.41)

OpenStudy (anonymous):

D ( -.59, -3.41)

OpenStudy (anonymous):

\[x _{1,2} =( -b \pm \sqrt{b^{2} - 4ac})/2a\] \[x_{1,2} = (-4 \pm \sqrt{16-8})/2\] \[x_{1,2} = -2 \pm \sqrt{2}\]